Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-15r + 12s = 24
30r – 24s = 48

Found 2 solutions by jorel1380, Alan3354:
Answer by jorel1380(3719)

-15r+12s=24
30r-24s=48
---------------------
-30r+24s=48
30r-24s=48
0=96
This system has no solution..

Answer by Alan3354(69443)

-15r + 12s = 24 times -2 --> 30r - 24s = -48
30r – 24s = 48
30r - 24s = -48
--------------- Subtract
0 = 96 --> no solutions
If you graph them, they're parallel. No intersection, no solution.